# A New One-Shot Federated Learning Framework for Medical Imaging Classification with Feature-Guided Rectified Flow and Knowledge Distillation

### v1

@@ Abstract
In multi-center scenarios, One-Shot Federated Learning (OSFL) has attracted increasing attention due to its low communication overhead, requiring only a single round of transmission. However, existing generative model-based OSFL methods suffer from low training efficiency and potential privacy leakage in the healthcare domain. Additionally, achieving convergence within a single round of model aggregation is challenging under non-Independent and Identically Distributed (non-IID) data. To address these challenges, in this paper a modified OSFL framework is proposed, in which a new Feature-Guided Rectified Flow Model (FG-RF) and Dual-Layer Knowledge Distillation (DLKD) aggregation method are developed. FG-RF on the client side accelerates generative modeling in medical imaging scenarios while preserving privacy by synthesizing feature-level images rather than pixel-level images. To handle non-IID distributions, DLKD enables the global student model to simultaneously mimic the output logits and align the intermediate-layer features of client-side teacher models during aggregation. Experimental results on three non-IID medical imaging datasets show that our new framework and method outperform multi-round federated learning approaches, achieving up to 21.73% improvement, and exceeds the baseline FedISCA by an average of 21.75%. Furthermore, our experiments demonstrate that feature-level synthetic images significantly reduce privacy leakage risks compared to pixel-level synthetic images.
@@ 1 Introduction
FL is a distributed machine learning framework in which multiple clients independently train local models and share only model updates with a central server for aggregation. This approach effectively protects user data privacy while enabling collaborative training. However, non-Independent and Identically Distributed (non-IID) data across clients remains a major challenge in FL [25, 27, 43], significantly affecting aggregation performance. Some studies have attempted to address the non-IID problem, such as FedProx [21], SCAFFOLD [15], and FedNova [34], focus on improving the efficiency and stability of model aggregation by addressing issues like client model drift, gradient inconsistency, and objective misalignment. Nevertheless, these methods rely solely on the raw data available on the client devices, without incorporating any generative models. The reliance on direct model updates and multiple rounds of communication to converge the global model creates significant communication overhead, limiting the scalability and applicability of FL in real-world scenarios.
@@
In OSFL with generative models, researchers have also adopted knowledge distillation (KD) [12] to facilitate global model convergence under non-IID data distributions, as relying solely on a single round of model aggregation can introduce bias, making convergence difficult (see Figure 1). In previous works [40, 14, 36], during the KD process, the global model (student model) learns not only from synthetic data but also from the output logits of the client models (teacher models). The logits from teacher models can capture diverse client data distributions, thus aiding aggregation under non-IID conditions. However, medical images exhibit greater feature differences and more complex decision boundaries compared to natural images [22, 14]. Moreover, potential distribution shifts in synthetic data may corrupt logit information from client models, thereby impairing aggregation performance. Thus, KD aggregation methods that rely solely on teacher logits are insufficient to ensure robust global model performance.
@@
To address the challenges of generative model-based OSFL for medical image classification, we propose a modified framework incorporating a new Feature-Guided Rectified Flow Model (FG-RF) and a Dual-Layer Knowledge Distillation (DLKD) method. Specifically, in the FG-RF model, a feature extractor is first trained to extract features from the original client images. The Rectified Flow Model (RFM) [23] then learns the velocity field at the feature level, achieving better privacy protection with shorter training and inference times compared to DDPM. To further enhance global model aggregation, the DLKD is designed to enable the global model (student model) to not only receive the output logits of client models (teacher models) but also to align its intermediate-layer features with those of the client models. This allows the student model to capture more knowledge like feature representations which cannot be fully transferred through final logits, helping to solve problems of feature difference in non-IID medical images and enhance the generalization ability of the global model. Moreover, the DLKD strategy helps mitigate performance degradation caused by the distribution gap between synthetic and original features.
@@
We design a new OSFL framework specialized for medical image classification incorporating diffusion model and knowledge distillation (KD). This approach uses KD to achieve single-round aggregation with synthetic data matching client distributions, avoids multiple communication rounds in traditional FL by requiring clients to upload only their trained models.
@@ 3 Methods
Our proposed framework for One-Shot Federating learning (OSFL) in medical imaging classification is mainly composed of two components: the Feature-Guided Rectified Flow Model (FG-RF) and Dual-Layer Knowledge Distillation (DLDK) method. In FG-RF, each client first trains a feature extractor to obtain features for subsequent Rectified Flow Model (RFM) training due to privacy concerns. The RFM is then trained based on the extracted features and subsequently transmitted to the server. Meanwhile, DLKD enables the global model to learn not only from the output logits of client models but also from dynamically aligned intermediate-layer features, enhancing generalization and mitigating overfitting.
@@ 3.1 Problem formulation
where is s composed of synthetic features generated from all clients’ RFMs, is the aggregated teacher model averaged from all client classifiers , is the cross-entropy loss used to train the student model for classification, and is the knowledge distillation loss.
@@ 3.2 Feature-Guided Rectified Flow Model
The feature extractor consists of two components: a convolutional layer and the tangent activation function (tanh). The convolutional layer is chosen because it not only effectively extracts image features but also irreversibly perturbs the original image, thereby preventing excessive information leakage. In our experiments, the selected convolutional layer typically preserve the number of channels, ensuring optimal compatibility with the input requirements of the subsequent RFM network training. The design choice of the tanh(x) activation function is motivated by the need to match the input format of RFM, as its output range of aligns with the image data range used in diffusion models. Moreover, tanh(x) compresses the original image features, further enhancing privacy protection. The combination of the convolutional layer and the tanh(x) activation function not only enhances privacy protection but also ensures a lightweight and simple implementation, making it easy to integrate into any classification network training pipeline.
@@ 3.3 Dual-Layer Knowledge Distillation
To train the server-side student model, we adopt Dual-Layer Knowledge Distillation (DLKD). Unlike prior works [40, 14, 36] that only distill the output logits from client models, DLKD additionally aligns intermediate-layer features between client and student models. Specifically, the intermediate layers of the teacher model contain hierarchical feature representations, which cannot be fully transferred through the final output alone. Aligning intermediate layers allows the student model to learn better feature representations, improving convergence speed and generalization ability. Figure 5 shows the t-SNE visualization of the TB dataset before and after DLKD training. As shown in Figure 5, the feature distributions are almost fully merged after DLKD training, implying that DLKD significantly reduces feature distribution discrepancies across clients.
@@ 3.4 Theoretical analysis
The analysis of the upper bound of DLKD error. Since our OSFL framework employs synthetic data training the final global model, we theoretically derive the upper bound of the generalization error of the global model on the real dataset. This analysis demonstrates that our proposed DLKD can effectively reduce the error introduced by generated data.
@@
Our DLKD reduce both terms. Unlike traditional distillation that only matches output logits, DLKD additionally aligns intermediate-layer features, enhancing both feature extraction and classification robustness. Feature alignment also maps generated features closer to the real data subspace, effectively reducing . As a result, both the distillation error and distributional gap are minimized, leading to better generalization on real-world data.
@@ 4.2 Baseline methods comparison
We compare our method with eleven baseline methods grouped into three categories: 1) centralized training and multi-round Federated Learning (FL): Centralized training serves as the performance upper bound for comparison. We select FedAvg [29], FedProx [21], FedNova [34], and MOON [20] as representatives of multi-round FL. In the experiments, we set 100 communication rounds for multi-round FL. 2) One-Shot Federated Learning (OSFL) based on adversarial approaches: In this category, we select DENSE [40] and FedISCA [14], both of which employ adversarial training to train the synthesizer. It is worth noting that FedISCA is specifically designed for medical image classification. 3) OSFL based on diffusion models: For this category, we compare our method with four approaches: FedCADO [36] uses client-trained classifiers to guide server-side diffusion model inference. FGL [41] transmits prompts generated by BLIP-2 to the pretrained diffusion model on the server. FedDEO [38] trains and sends text descriptions to the pretrained diffusion models on the server. The DDPM baseline generates feature-level images using a DDPM model, with all other operations consistent with our method, serving as a benchmark for generative model quality. We evaluate the classification accuracy of our proposed method against the baseline methods, as shown in Table 1. The experiment details can be found in Appendix A [28].
@@
FedCADO | FGL | FedDEO | DDPM |
Brain Tumor | ![]() |
Chest X-Ray | ![]() |
HAM10000 | ![]() |
Table 2 compares the image features generated by our method with the original images and baseline methods (see more generated results of our method in Appendix B [28]). Our method, DENSE, and FedISCA generate image features, while the others synthesize pixel-level images. The results show that DENSE and FedISCA struggle to produce high-resolution features, resulting in detail loss and impaired model aggregation. FedCADO, FGL, and FedDEO generate synthetic data with large distribution shifts, also degrading aggregation performance. In contrast, our method generates high-quality features with slight blurring, effectively balancing utility and privacy.
@@
To demonstrate that the efficiency of FG-RF, we compare the training and inference times of DDPM and FG-RF, as shown in Table 3. The results show that FG-RF significantly reduces both training and inference times across all three datasets, indicating that the RFM backbone is more suitable for OSFL scenarios.

### v2

@@ Abstract
In multi-center scenarios, One-Shot Federated Learning (OSFL) has attracted increasing attention due to its low communication overhead, requiring only a single round of transmission. However, existing generative model-based OSFL methods suffer from low training efficiency and potential privacy leakage in the healthcare domain. Additionally, achieving convergence within a single round of model aggregation is challenging under non-Independent and Identically Distributed (non-IID) data. To address these challenges, in this paper a modified OSFL framework is proposed, in which a new Feature-Guided Rectified Flow Model (FG-RF) and Dual-Layer Knowledge Distillation (DLKD) aggregation method are developed. FG-RF on the client side accelerates generative modeling in medical imaging scenarios while preserving privacy by synthesizing feature-level images rather than pixel-level images. To handle non-IID distributions, DLKD enables the global student model to simultaneously mimic the output logits and align the intermediate-layer features of client-side teacher models during aggregation. Experimental results on three non-IID medical imaging datasets show that our new framework and method outperform multi-round federated learning approaches, achieving up to 21.73% improvement, and exceed the baseline FedISCA by an average of 21.75%. Furthermore, our experiments demonstrate that feature-level synthetic images significantly reduce privacy leakage risks compared to pixel-level synthetic images. The code is available at https://github.com/LMIAPC/one-shot-fl-medical.
@@ 1 Introduction
FL is a distributed machine learning framework in which multiple clients independently train local models and share only model updates with a central server for aggregation. This approach effectively protects user data privacy while enabling collaborative training. However, non-Independent and Identically Distributed (non-IID) data across clients remains a major challenge in FL [25, 27, 43], significantly affecting aggregation performance. Some studies have attempted to address the non-IID problem, such as FedProx [21], SCAFFOLD [15], and FedNova [34], which focus on improving the efficiency and stability of model aggregation by addressing issues like client model drift, gradient inconsistency, and objective misalignment. Nevertheless, these methods rely solely on the raw data available on the client devices, without incorporating any generative models. The reliance on direct model updates and multiple rounds of communication to converge the global model creates significant communication overhead, limiting the scalability and applicability of FL in real-world scenarios.
@@
In OSFL with generative models, researchers have also adopted knowledge distillation (KD) [12] to facilitate global model convergence under non-IID data, as relying solely on a single round of model aggregation can introduce bias, making convergence difficult (see Figure 1). In previous works [40, 14, 36], during the KD process, the global model (student model) learns not only from synthetic data but also from the output logits of the client models (teacher models). The logits from teacher models can capture diverse client data distributions, thus aiding aggregation under non-IID conditions. However, medical images exhibit greater feature differences and more complex decision boundaries compared to natural images [22, 14]. Moreover, potential distribution shifts in synthetic data may corrupt logit information from client models, thereby impairing aggregation performance. Thus, KD aggregation methods that rely solely on teacher logits are insufficient to ensure robust global model performance.
@@
To address the challenges of generative model-based OSFL for medical image classification, we propose a modified framework incorporating a new Feature-Guided Rectified Flow Model (FG-RF) and a Dual-Layer Knowledge Distillation (DLKD) method. Specifically, in the FG-RF model, a feature extractor is first trained to extract features from the original client images. The Rectified Flow Model (RFM) [23] then learns the velocity field at the feature level, achieving better privacy protection with shorter training and inference times compared to DDPM. To further enhance global model aggregation, the DLKD is designed to enable the global model (student model) to not only receive the output logits of client models (teacher models) but also to align its intermediate-layer features with those of the client models. This allows the student model to capture more knowledge like feature representations, which cannot be fully transferred through final logits, helping to solve problems of feature difference in non-IID medical images and enhance the generalization ability of the global model. Moreover, the DLKD strategy helps mitigate performance degradation caused by the distribution gap between synthetic and original features.
@@
We design a new OSFL framework specialized for medical image classification, incorporating diffusion model and knowledge distillation (KD). This approach uses KD to achieve single-round aggregation with synthetic data matching client distributions. It avoids multiple communication rounds in traditional FL by requiring clients to upload only their trained models.
@@ 3 Methods
Our proposed framework for One-Shot Federated Learning (OSFL) in medical imaging classification is mainly composed of two components: the Feature-Guided Rectified Flow Model (FG-RF) and Dual-Layer Knowledge Distillation (DLDK) method. In FG-RF, each client first trains a feature extractor to obtain features for subsequent Rectified Flow Model (RFM) training due to privacy concerns. The RFM is then trained based on the extracted features and subsequently transmitted to the server. Meanwhile, DLKD enables the global model to learn not only from the output logits of client models but also from dynamically aligned intermediate-layer features, enhancing generalization and mitigating overfitting.
@@ 3.1 Problem formulation
where is composed of synthetic features generated from all clients’ RFMs, is the aggregated teacher model averaged from all client classifiers , is the cross-entropy loss used to train the student model for classification, and is the knowledge distillation loss.
@@ 3.2 Feature-Guided Rectified Flow Model
The feature extractor consists of two components: a convolutional layer and the tangent activation function (tanh). The convolutional layer is chosen because it not only effectively extracts image features but also irreversibly perturbs the original image, thereby preventing excessive information leakage. In our experiments, the selected convolutional layer typically preserves the number of channels, ensuring optimal compatibility with the input requirements of the subsequent RFM network training. The design choice of the tanh(x) activation function is motivated by the need to match the input format of RFM, as its output range of aligns with the image data range used in diffusion models. Moreover, tanh(x) compresses the original image features, further enhancing privacy protection. The combination of the convolutional layer and the tanh(x) activation function not only enhances privacy protection but also ensures a lightweight and simple implementation, making it easy to integrate into any classification network training pipeline.
@@ 3.3 Dual-Layer Knowledge Distillation
To train the server-side student model, we employ Dual-Layer Knowledge Distillation (DLKD). Unlike prior works [40, 14, 36] that only distill the output logits from client models, DLKD additionally aligns intermediate-layer features between client and student models. Specifically, the intermediate layers of the teacher model contain hierarchical feature representations, which cannot be fully transferred through the final output alone. Aligning intermediate layers allows the student model to learn better feature representations, improving convergence speed and generalization ability. Figure 5 shows the t-SNE visualization of the TB dataset before and after DLKD training. As shown in Figure 5, the feature distributions are almost fully merged after DLKD training, implying that DLKD significantly reduces feature distribution discrepancies across clients.
@@ 3.4 Theoretical analysis
The analysis of the upper bound of DLKD error. Since our OSFL framework employs synthetic data for training the final global model, we theoretically derive the upper bound of the generalization error of the global model on the real dataset. This analysis demonstrates that our proposed DLKD can effectively reduce the error introduced by generated data.
@@
Our DLKD reduces both terms. Unlike traditional distillation that only matches output logits, DLKD additionally aligns intermediate-layer features, enhancing both feature extraction and classification robustness. Feature alignment also maps generated features closer to the real data subspace, effectively reducing . As a result, both the distillation error and distributional gap are minimized, leading to better generalization on real-world data.
@@ 4.2 Baseline methods comparison
We compare our method with eleven baseline methods grouped into three categories: 1) centralized training and multi-round Federated Learning (FL): Centralized training serves as the performance upper bound for comparison. We select FedAvg [29], FedProx [21], FedNova [34], and MOON [20] as representatives of multi-round FL. In the experiments, we set 100 communication rounds for multi-round FL. 2) One-Shot Federated Learning (OSFL) based on adversarial approaches: In this category, we select DENSE [40] and FedISCA [14], both of which employ adversarial training to train the synthesizer. It is worth noting that FedISCA is specifically designed for medical image classification. 3) OSFL based on diffusion models: For this category, we compare our method with four approaches: FedCADO [36] uses client-trained classifiers to guide server-side diffusion model inference. FGL [41] transmits prompts generated by BLIP-2 to the pretrained diffusion model on the server. FedDEO [38] trains and sends text descriptions to the pretrained diffusion models on the server. The DDPM baseline generates feature-level images using a DDPM model, with all other operations consistent with our method, serving as a benchmark for generative model quality. We evaluate the classification accuracy of our proposed method against the baseline methods, as shown in Table 1. The experimental details can be found in the supplementary material [28].
@@
FedCADO | FGL | FedDEO |
| Brain Tumor | |||||||||||||||||||
| Chest X-Ray | |||||||||||||||||||
| HAM10000 |
Table 2 compares the image features generated by our method with the original images and baseline methods (See supplementary material for more results [28]). DENSE, FedISCA, DDPM and our method generate image features, while the others synthesize pixel-level images. The results show that DENSE and FedISCA struggle to produce high-resolution features, resulting in detail loss and impaired model aggregation. FedCADO, FGL, and FedDEO generate synthetic data with large distribution shifts, also degrading aggregation performance. In contrast, our method generates high-quality features with slight blurring, effectively balancing utility and privacy.
@@
To demonstrate the efficiency of FG-RF, we compare the training and inference times of DDPM and FG-RF, as shown in Table 3. The results show that FG-RF significantly reduces both training and inference times across all three datasets, indicating that the RFM backbone is more suitable for OSFL scenarios.